Solve for X and Y, two system equation:
Equation1 4X+Y=17 Equation2 2X+3Y=21

Answer by checkley77(12844)

4X+Y=17 MULTIPLY BY -3 & ADD.
2X+3Y=21
-12X-3Y=-51
--------------
-10X=-30
X=-30/-10
X=3 ANS.
4*3+Y=17
12+Y=17
Y=17-12
Y=5 ANS.
PROOF:
2*3+3*5=21
6+15=21
21=21